OpenStudy (anonymous):

Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. please show work. 2x + 5 + 4x = 7x + 5 – x

OpenStudy (anonymous):

2x+4x=6x+5 7x-x+5=6x+5 so you have 6x-6x=5-5 you answer is 0